94. Marginal revenue. A defense attorney charges her client $4,000 plus $120 per hour. The formula R = 120n + 4000 gives her revenue in dollars for n hours of work. What is her revenue for 100 hours of work? What is her revenue for 101 hours of work? By how much did the one extra hour of work increase the revenue? (The increase in revenue is called the marginal revenue for the 101st hour.)

Answer by elima(1433)

R = 120n + 4000
For 100 hours of work?
R=120(100)+4000=12000+4000=16000
For 101 hours of work?
R=120(101)+4000 = 12120+4000 = 16120
Increase = 120
:)